### v1

# GenSwarm: Scalable Multi-Robot Code-Policy Generation and Deployment via Language Models

Abstract
The development of control policies for multi-robot systems traditionally follows a complex and labor-intensive process, often lacking the flexibility to adapt to dynamic tasks. This has motivated research on methods to automatically create control policies. However, these methods require iterative processes of manually crafting and refining objective functions, thereby prolonging the development cycle. This work introduces GenSwarm, an end-to-end system that leverages large language models to automatically generate and deploy control policies for multi-robot tasks based on simple user instructions in natural language. As a multi-language-agent system, GenSwarm achieves zero-shot learning, enabling rapid adaptation to altered or unseen tasks. The white-box nature of the code policies ensures strong reproducibility and interpretability. With its scalable software and hardware architectures, GenSwarm supports efficient policy deployment on both simulated and real-world multi-robot systems, realizing an instruction-to-execution end-to-end functionality that could prove valuable for robotics specialists and non-specialists alike.The code of the proposed GenSwarm system is available online: https://github.com/WindyLab/GenSwarm.
Introduction
Multi-robot systems show significant promise for applications both indoors (for example, factory floors, warehouses, hospitals) and outdoors (for example, transport, inspection, farming, disaster response) [1, 2]. The present paradigm of developing multi-robot systems follows a complex and labor-intensive process that involves steps like task analysis, algorithm design, code programming, simulation validation, and real-world deployment. This paradigm requires skilled professionals who are familiar with both theories and software/hardware implementation, incurring high costs in human resources. Moreover, it does not adapt well to dynamically changing tasks: the emergence of a new task requires the repetition of the complex process.
Automatic generation and deployment of control policies for multi-robot systems is an appealing paradigm, as it promises substantial savings in terms of human effort and other resources [3, 4, 5]. However, this paradigm is nontrivial to realize as a multi-robot system as a whole cannot be programmed directly; rather, a desired collective behavior can be achieved only by programming each individual robot, which relies on its locally available information. Previous methods for automatic development of multi-robot swarming are primarily based on optimization techniques [3, 5]. For instance, an objective function is first crafted to mathematically describe a desired task and then optimized to generate policies through methods such as evolutionary computation [6, 7, 5] or systematic search [8]. Despite their promise, these optimization methods face the common limitation of requiring manual crafting of objective functions.
Recent advances in large language models (LLMs) [9, 10] and vision language models (VLMs) [11, 12] offer new paradigms for developing robotic systems. In one paradigm, a language model can be deployed onboard a robot to directly make decisions online [13, 14]. Due to the generality of language models, this paradigm could be used to address open-ended tasks [15, 16, 17]. However, it faces challenges in terms of reproducibility, interpretability, and hallucination. In another paradigm, a language model is used to generate executable code policies that are subsequently uploaded for execution on-board robots. A representative method that falls into this paradigm is Code-as-Policy [18, 19, 20]. Due to the white-box nature of executable code, this paradigm offers high reproducibility and interpretability. Moreover, since executable code usually requires fewer resources than LLMs, this paradigm also enables real-time control on low-cost robot platforms. This is especially relevant for large-scale swarming systems, where collective behaviors emerge from the interactions among robots with exceedingly limited onboard resources [21, 22, 23, 24, 25, 26]. Therefore, this code-policy paradigm is adopted in the present work.
Despite the promise of the code-policy paradigm, current studies mainly focus on single-robot systems [18, 27, 28]. For multi-robot systems, the development of control policies faces additional challenges. First, the design of policies must consider a robot’s interactions with its peers. In some situations the robot may compete with its peers, for example, for limited resources, whereas in others it may cooperate with its peers to achieve a common goal [26, 29, 5]. Second, the deployment and maintenance of policies require scalable software and hardware systems, which is particularly relevant for robotic swarms that may have a large number of robots. Third, to maximize the utility of a multi-robot system, it needs to support a wide range of tasks. Although some studies used LLMs to realize specific tasks, such as cooperative navigation [30], flocking [31], dancing [32, 33], and manipulation [34], the generality of the proposed methods, beyond the specific use case, is yet to be demonstrated. A few further studies have explored LLMs with groups of agents [35, 36, 37], but the agents therein are simulated non-embodied agents; hence the findings may be difficult to translate to real multi-robot systems.
Here, we propose GenSwarm, an end-to-end system that can automatically generate and deploy code policies for versatile multi-robot tasks. GenSwarm enables users to program a group of robots using simple natural language instructions. The user instructions are automatically processed via a pipeline of components, including constraint analysis, policy design, policy generation, policy deployment in simulation environments, policy deployment on real-world robots, and policy improvement based on feedback. These components are respectively empowered by LLM agents. Therefore, GenSwarm realizes generative multi-robot swarming via the cooperation among multiple language agents.
GenSwarm can deploy the generated code policies as well as the required runtime environments fully automatically on simulated and real-world robots, thus achieving true end-to-end functionality. The deployment is facilitated by a scalable multi-robot platform that features novel software and hardware architectures. GenSwarm enables zero-shot policy generation without the need for context learning based on demonstrative examples. When altered or unseen tasks arise, the system can re-generate and re-deploy policies in response to user requests, thereby offering high adaptability for dynamic tasks. Furthermore, due to the use of code policies, the approach is not only reproducible and interpretable, but also suitable for real-time execution on robots with limited onboard resources.
Overview of GenSwarm
The pipeline of GenSwarm consists of three modules: task analysis, code generation, and code deployment and improvement (Fig. 1).
Fig. 1: The pipeline of GenSwarm. GenSwarm consists of three modules: task analysis, code generation, and code deployment and improvement. The task analysis module extracts constraints from user instructions and builds a skill library. The code generation module uses a skill graph to hierarchically create and refine Python functions, ensuring constraint alignment and code reusability. Finally, the code deployment and improvement module enables automatic code deployment in simulation and real-world platforms, incorporating feedback from video analysis and human input to refine policies.
The task analysis module takes as input user instructions in the form of natural language about the desired multi-robot task (Fig. 1a). For example, to achieve a predator-prey encircling task, the user instruction could be “The robots need to surround the target prey by evenly distributing themselves along a circle with a radius of 1, centered on the prey.” From the user instruction, an LLM agent extracts constraints that compose a constraint pool. Each constraint specifies what a robot shall or shall not do, such as reaching a target location or not colliding with obstacles. Since the constraint pool comprehensively captures the task requirements, every subsequent step must align with the constraints, thereby ensuring the task is achieved as intended. Based on the constraints, an LLM agent generates a skill library where each skill corresponds to a Python function. At this stage, merely the function’s name and description are generated; the main body of the function will be generated at a later stage. Skills can be classified as either global or local. Global skills involve global coordination such as goal assignment, whereas local skills are executed onboard each robot based on locally available information.
The code generation module generates the code for the main body of each skill function (Fig. 1b). First, a skill graph is constructed by an LLM agent to describe the hierarchical dependencies between the skills and to indicate the constraints that each skill must satisfy. The skill graph guides the code generation process: low-level skills are generated first, and high-level skills thereafter, thereby enhancing code reuse and reducing the need for repetitive code modifications due to lower-level errors. Once the main body of each skill function has been generated, an LLM agent reviews whether the function aligns with the associated constraints, and makes modifications if necessary. After the review, static code checks are performed, and an LLM agent makes modifications if necessary, ensuring the code is executable.
The code deployment and improvement module realizes automatic code deployment in simulated and real-world robotic platforms (Fig. 1c). It relies on novel hardware and software systems, which will be detailed in the following section. It introduces multi-modal feedback mechanisms that can automatically identify issues during execution and effectively adjust policies based on feedback. Specifically, execution results in the simulation can be automatically collected in the format of video clips. A VLM agent assesses the video clips to generate feedback on whether the desired task is successfully completed. In addition, an interface for human feedback is incorporated to enable users to efficiently modify the policy by providing natural language feedback.
Software and Hardware Platform
Fig. 2: Software components of GenSwarm. A control station generates the required code based on the proposed pipeline and uses Ansible to wirelessly connect to each robot. First, each robot runs Playbook-defined tasks, such as installing and configuring the Docker environment. Then, two pre-built Docker images are pulled: one with the ROS environment for robot operation, and the other with the Python environment for code execution. Once the environments are ready, the generated code is transmitted to all robots and then executed onboard.
Fig. 3: Hardware components of GenSwarm. As a major upgrade of our previous robotic platform [26], each robot has the onboard computational, control, and communication resources to support autonomous code deployment and execution. The multi-robot system features one-click all start, one-click all sleep, and wireless data retrieval functions that can significantly reduce experimental costs. Since the robots do not have onboard vision systems, the perception was emulated with relevant motion information being collected by an indoor motion capture system, and then distributed to the robots through an MQTT coordination server, ensuring each robot receives only the local information of its surroundings.
Automatic deployment is nontrivial as code execution depends on complex runtime environments consisting of various software packages. Manually installing and configuring the runtime environments on each robot would be inefficient as the time required scales linearly with the number of robots; for example, deploying on 10 robots would take 10 times longer than deploying on one robot. This would make deployment on large-scale swarms impractical.
GenSwarm possesses a software framework that can automatically deploy both the runtime environments and the generated code across all the robots in near constant time regardless of the number of robots. In our experiments, automatically deploying the runtime environments on all the robots takes about two minutes, whereas automatically deploying the generated code takes mere seconds. This makes the system particularly well-suited for large-scale swarming systems, where consistent and rapid deployment is essential.
The software framework is illustrated in Fig. 2. A control station first generates the required code based on the pipeline described earlier and connects with each robot through Ansible via WiFi and SSH (Methods). With predefined automated scripts in the format of Playbook, each robot performs a series of tasks such as installing and configuring the Docker environment. After the Docker environment is ready, two pre-built Docker images are pulled: one containing the ROS environment used for robot operation, and the other containing the Python environment required for code execution. Once the execution environments are ready, the generated code is transmitted to all robots and then executed onboard. The proposed software framework heavily relies on two techniques, Ansible and Docker (Methods), which work together to simplify and streamline the code deployment on multiple robots. This integration ensures that the deployment process is both repeatable and efficient, drastically reducing the time required to make a group of robots operational. In addition, the framework is realized by a modular architecture, whose details are given in Methods.
The hardware framework is illustrated in Fig. 2. A new multi-robot platform, which is a major upgrade of our previous robotic platform [26], was developed to support GenSwarm. Each ground robot has onboard computational, control, and communication resources that are necessary for autonomous code deployment and execution [38]. Considering that multi-robot experiments involve a large number of operations, such as starting and shutting down robots, we developed novel features for the multi-robot platform such as one-click all start, one-click all sleep, and wireless data retrieval, significantly reducing experimental costs.
It is worth mentioning that since the robots do not possess onboard vision systems, the perception is emulated in a distributed manner. Specifically, relevant motion information of all the robots is collected by an indoor positioning system first and then distributed to all the robots through an MQTT coordination server so that each robot receives only information about its surroundings (Fig. 3). In the future, this coordination server could be replaced with on-board vision systems. Another merit of the MQTT coordination server is to enable large-scale hardware-in-the-loop swarming simulation and seamless transition from simulation to real-world implementation [38].
Demonstration of GenSwarm
Fig. 4: A demonstration of the complete workflow of GenSwarm. a. The user specified a predator-prey encircling task via natural language. b. Six constraints were extracted from the task by LLM agents. c. Six skills were generated and categorized into global (e.g., goal assignment) and local (e.g., update velocity). d. Based on those skills and a consequently generated skill graph, LLM agents generated and reviewed the main-body code of each skill function. e. The code was deployed in simulation environments, reviewed via video feedback by VLM agents, and refined with human feedback. Once validated, it was automatically deployed on real-world robots.
To demonstrate the workflow of GenSwarm, we show the end-to-end generation process of a representative multi-robot task “predator-prey encircling” (Fig. 4). In this task, multiple predator robots should follow and surround a prey robot that moves randomly. The user instruction is shown in Fig. 4a. From the user instruction, six constraints were generated by an LLM agent (Fig. 4b). For instance, one of the constraints is “CollisionAvoidance” with the description as “Ensure a minimum distance greater than the sum of the robot’s radius, other robots’ radii, and a pre-
defined distance_threshold from all other robots and obstacles within the perception
range
Based on those constraints, six skills (merely the names and descriptions) are generated (Fig. 4c). One of them is a global skill that will be executed on the control station, whereas the others are local skills that will be executed on each robot in a distributed manner. The purpose of the global skill, named “Allocate_initial_angles”, is goal assignment, that is, to assign the desired relative angular position of each robot when encircling the target. Goal assignment is a common technique adopted in multi-robot tasks, especially when there is a global constraint such as a geometric shape that multiple robots must satisfy [26, 39].
Based on those skills and the consequently generated skill graph that describes their hierarchical dependencies, LLM agents further generate and then review the main-body code of each skill function (Fig. 4d). Logical or grammatical code errors can be identified and corrected. Once the skill functions pass the review process and static code analysis, they are automatically deployed and executed in the simulation environment. Then, a VLM agent reviews the video clip of the simulation execution and provides feedback for improving the code (Fig. 4e). After that, the generated code is automatically deployed on real-world robotic platforms. It is worth mentioning that human feedback can be incorporated to adjust the code policy (Fig. 4e). For instance, if the human feedback is “Forming a circle with a radius of 1 meter seems a bit large; perhaps 0.8 meters would be better.”, GenSwarm can adjust the corresponding parameter from 1 to 0.8, enabling efficient human-in-the-loop policy adjustment.
Non-stop one-take videos are attached to show the complete workflows of GenSwarm (Movies 1 and 2). In terms of time consumption, the steps of code generation, deployment onto real-world robots, and improvement based on human feedback took approximately six, two, and two minutes, respectively. The time duration of code generation can be significantly shortened if LLMs’ efficiency can be improved in the future. The time of deployment can be shortened to a few seconds if the runtime environment has been pre-installed on the robots and merely the generated code needs to be deployed.
As elaborated above, GenSwarm consists of multiple LLM agents that play different roles. Each LLM agent is set up in advance by a prompt involving role description, environment description, robot description, and available APIs. For instance, regarding environment description, the prompt may be “The environment is composed of a 2D plane with obstacles and robots”. Regarding robot description, the prompt may be “The maximum speed of each agent is 0.5 m/s”. Regarding APIs, the prompt may be “There are two types of APIs: local and global. Local APIs can only be called by the robot itself, and global APIs can be called by a cen-
tralized controller
”. Examples of local APIs are “get_self_position” and “get_surrounding_
robots_info
”. Examples of global APIs are “get_all_robots_id” and “get_all_robots_initial_
position
”. More information can be found in the code on GitHub (Code Availability).
Fig. 5: Results by GenSwarm for ten multi-robot tasks. The ten tasks include aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering. These tasks cover a wide range of scenarios, from cooperative to competitive, aiming to comprehensively evaluate the effectiveness of GenSwarm.
Fig. 6: Evaluation results. a. The success rates of GenSwarm across ten multi-robot tasks. One hundred independent trials, from user instructions to code execution in simulation, were run for each task. Hence, 1,000 trials in total were run and the average success rate was 81%. b. The comparison between GenSwarm, CaP, MetaGPT, and GenSwarm without VLM feedback across six representative tasks. One hundred independent trials, from user instructions to code execution in simulation, were run for each method and each task. The average success rates of GenSwarm, GenSwarm without VLM, CaP, and MetaGPT were 74%, 71%, 40%, and 31%, respectively.
Performance Evaluation
The performance of GenSwarm was evaluated on ten multi-robot tasks, including aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering (Fig. 5). These tasks cover a wide range of scenarios, from cooperative to competitive, aiming to comprehensively evaluate the effectiveness of GenSwarm. Details of the tasks and the evaluation metrics are given in Methods. The LLM used here was o1-mini, one of the state-of-the-art LLMs. One hundred independent trials, starting from user instruction to code execution in simulation, were run for each of the ten tasks. The average success rate over the 1,000 trials for 10 tasks was 81%. The respective success rate for each task is presented in Fig. 6a.
The performance of GenSwarm is also compared to two state-of-the-art methods, MetaGPT [35] and Code-as-Policy (CaP) [18]. MetaGPT and CaP are two seminal methods that represent the mainstream paradigms in the fields of complex software generation and robot policy generation, respectively. GenSwarm without VLM feedback was also evaluated. One hundred independent trials, from user instructions to code execution in simulation, were run for each method and each task. Six representative tasks were selected, and hence 1,800 trials in total were run. The LLM used here was GPT-4o. The comparison results are shown in Fig. 6b. As can be seen, GenSwarm achieved the highest average success rate, which was 74%, across different tasks. The average success rates of GenSwarm without VLM, CaP, and MetaGPT were 71%, 40%, and 31%, respectively. It is worth mentioning that to avoid underestimating the capabilities of CaP and MetaGPT, we discarded trials that were not executable due to code bugs and only counted the executable trials for CaP and MetaGPT. Nevertheless, the two methods still show lower performance than GenSwarm. By comparing Figs. 6a and 6b, it can be seen that different LLMs (o1-mini and GPT-4o) lead to similar success rates though there are subtle variations.
The user instructions have a significant impact on the success rate of GenSwarm (Extended Data Fig. 1). The success rate tends to be higher when comprehensive instructions are provided. When the instructions are ambiguous, the generated policies are more likely to encounter issues. Note that due to the limited reasoning ability and hallucinations of LLMs, even detailed instructions cannot guarantee a 100% success rate (Extended Data Fig. 1a). Moreover, it is still possible, albeit unlikely, to successfully accomplish the task when the instructions are ambiguous (Extended Data Fig. 1b-c).
Discussion
This work introduced GenSwarm, an end-to-end system that automatically generates and deploys code policies for versatile multi-robot tasks. As a significant step toward end-to-end generation, GenSwarm presents a novel paradigm that could potentially disrupt the current development process of multi-robot systems. However, GenSwarm has some limitations that could be addressed in the future. First, this study focussed on decision-making and control. Aspects such as sensing and navigation, which are important for practical applications, have not been incorporated. Developing and integrating onboard sensing into the system would be a valuable direction for future research. Second, although GenSwarm demonstrated high success rates for a variety of tasks, there is still potential to improve the success rate and address more complex multi-robot tasks. We believe that achieving these improvements might be challenging when relying solely on LLMs. Therefore, combining language models with other techniques, such as multi-agent reinforcement learning, could be a promising approach. Compared to language models, reinforcement learning is better suited for generating more sophisticated and optimal policies, making it a valuable complement to GenSwarm.
Method
Ten Swarming Tasks
The ten swarming tasks considered in this work are aggregation, flocking, shaping, encircling, crossing, coverage, exploration, pursuing, bridging, and clustering. The following gives the user instructions and evaluation metrics of each task. Multiple metrics may be used to evaluate a task from different aspects. It must be noted that none of these metrics were used when creating code policies using GenSwarm; they are merely used for the purpose of post-evaluation. A task is regarded as successful when all of its corresponding metrics exceed certain predefined thresholds. In this way, we can automatically calculate the success rate of each task. The termination of a simulation trial is triggered when the execution time exceeds certain values or the task has finished in the sense that, for example, all the robots succeed in reaching their desired positions.
  • Aggregation task
    User instruction:The robots need to aggregate as quickly as possible and avoid
    colliding with each other.
    Evaluation metric: Maximum of minimum distances, denoted as $d_{\text{maxmin}}$: It quantifies the largest minimum distance between each robot and its closest neighbor. It is defined as
    $d_{\text{maxmin}}=\max_{i}\min_{j\neq i}||\mathbf{p}_{i}-\mathbf{p}_{j}||$
    where $||\mathbf{p}_{i}-\mathbf{p}_{j}||$ is the Euclidean distance between robots $i$ and $j$. The task is regarded as successful if the value of this metric is less than 1.
  • Flocking task
    User instruction:The robots must form a cohesive flock, cooperating with all
    others in the environment. The three main behaviors are cohesion, alignment, and
    separation: cohesion maintains connectivity, alignment ensures synchronized move-
    ment, and separation prevents collisions by keeping robots at least 0.5 meters
    apart.
    Evaluation metrics: The flocking task is evaluated based on two metrics. The task is treated as successful when both metrics exceed their corresponding thresholds.
    1) Spatial Variance, denoted as $\text{Var}_{\text{spat}}$: It quantifies how spread out the robots are. It is defined as
    $\text{Var}_{\text{spat}}=\sum_{d\in\{x,y\}}\text{Var}(P_{d})$
    where $\text{Var}(P_{d})$ is the variance of the robot positions along the $d$ dimension (either $x$ or $y$). The task is regarded as successful if the value of this metric is less than 1.
    2) Mean Dynamic Time Warping (DTW) Distance, denoted as $d_{\text{DTW}}$: This metric quantifies the similarity between the trajectories of all robots. It is defined as
    $d_{\text{DTW}}=\frac{1}{M}\sum_{i<j}\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})$
    where $M$ is the total number of robot pairs, $\mathbf{T}_{i}$ is the trajectory of robot $i$, and $\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})$ is the DTW distance between $\mathbf{T}_{i}$ and $\mathbf{T}_{j}$. Here, $\mathbf{T}_{i}=\{(x_{i}^{1},y_{i}^{1}),\dots,(x_{i}^{1,000},y_{i}^{1,000})\}$ and $\mathbf{T}_{j}=\{(x_{j}^{1},y_{j}^{1}),\dots,(x_{j}^{1,000},y_{j}^{1,000})\}$. The DTW distance between them is defined as [40]
    $\text{DTW}(\mathbf{T}_{i},\mathbf{T}_{j})=\min_{W\in\Omega}\sum_{(a,b)\in W}d\bigl{(}(x_{i}^{a},y_{i}^{a}),(x_{j}^{b},y_{j}^{b})\bigr{)},$
    where $W$ is the warping path, a valid alignment between $\mathbf{T}_{i}$ and $\mathbf{T}_{j}$ that satisfies constraints such as boundary, continuity, and monotonicity. The function $d(\cdot,\cdot)$ is the Euclidean distance.
    The task is regarded as successful if the value of this metric is less than 500. Since each trajectory has 1,000 points, the threshold of 500 indicates that the average distance between pairs of points across two trajectories is less than 0.5.
  • Shaping task
    User instruction:The robots need to form a specific shape, with each robot assigned a unique point on that shape. The task requires each robot to move towards and
    maintain its assigned position on the target shape.
    Evaluation metric: Procrustes Distance, denoted as $d_{\text{proc}}$: It quantifies the similarity between the robot positions and the target shape. It is defined as
    $d_{\text{proc}}=\min_{\mathbf{Q}}\frac{1}{N}\sum_{i=1}^{N}||\mathbf{p}_{i}-\mathbf{Q}\mathbf{p}_{i,\text{T}}||^{2}$
    where $N$ is the total number of robots, $\mathbf{p}_{i}$ is the current position of robot $i$, $\mathbf{p}_{i,\text{T}}$ is the target position for robot $i$ on the straight line, and $\mathbf{Q}$ is the optimal permutation matrix to be solved. The task is regarded as successful if the value of this metric is less than 0.1.
  • Encircling task
    User instruction:The robots need to surround the target prey by evenly distributing themselves along a circle with a radius of 1, centered on the prey. Each robot is
    assigned a specific angle, and they must adjust their positions in real-time based
    on the prey’s movement to achieve coordinated encirclement.
    Evaluation metric: Mean distance error, denoted as $d_{\text{error}}$: It quantifies the average deviation of the robots’ distances from the desired radius. It is defined as
    $d_{\text{error}}=\frac{1}{N}\sum_{i=1}^{N}\big{|}||\mathbf{p}_{i}-\mathbf{p}_{\text{prey}}||-r_{\text{desired}}\big{|}$
    where $N$ is the total number of robots, $\mathbf{p}_{i}$ is the position of robot $i$, $\mathbf{p}_{\text{prey}}$ is the position of the prey, and $r_{\text{desired}}$ is the desired radius. The task is regarded as successful if the value of this metric is less than 0.1.
  • Crossing task
    User instruction:Each robot must maintain a distance of at least fifteen centi-
    meters from other robots and obstacles to avoid collisions while moving to the
    target point, which is the position of the robot that was farthest from it at the
    initial moment.
    Evaluation metric: Target Reach Ratio, denoted as $\rho_{\text{reach}}$: It quantifies the proportion of robots that successfully reached their target positions within a certain tolerance distance (typically 0.1 meters). It is defined as
    $\rho_{\text{reach}}=\frac{N_{\text{reach}}}{N}$
    where $N_{\text{reach}}$ is the number of robots reached targets. The task is regarded as successful if the value of this metric is equal to 1.
  • Coverage task
    User instruction:Divide the environment into sections equal to the number of robots. Each robot needs to move to the center of its assigned section to achieve full
    coverage of the environment.
    Evaluation metrics: The coverage task is evaluated based on two metrics. The task is treated as successful when both metrics exceed their corresponding thresholds.
    1) Area Ratio, denoted as $\rho_{\text{area}}$: It quantifies how much of the total area is occupied by the robots. It is defined as
    $\rho_{\text{area}}=\frac{A_{\text{occupied}}}{A_{\text{total}}}$
    where $A_{\text{occupied}}$ is the area occupied by the robots and $A_{\text{total}}$ is the total available area. More specifically, $A_{\text{occupied}}$ is calculated as $A_{\text{occupied}}=(x_{\max}-x_{\min})\times(y_{\max}-y_{\min})$, where $x_{\max}$ and $x_{\min}$ are the maximum and minimum x-coordinates among all the robots, respectively. The task is regarded as successful if the value of this metric is greater than 0.8.
    2) Variance of Nearest Neighbor Distances, denoted as $\text{Var}_{\text{NND}}$: It quantifies how evenly spaced the robots are from their nearest neighbors. It is defined as
    $\text{Var}_{\text{NND}}=\text{Var}(d_{\text{near}})$
    where $d_{\text{near}}$ is the Euclidean distance between each robot and its nearest neighbor. The task is regarded as successful if the value of this metric is less than 0.1.
  • Exploration task
    User instruction:The robots need to explore all the unknown areas. You are required to assign an optimal sequence of exploration areas to each robot based on the number of robots and the unexplored regions, and then the robots will gradually explore
    these areas.
    Evaluation metric: Landmark Visit Ratio, denoted as $\rho_{\text{visit}}$: It quantifies the proportion of unexplored areas (landmarks) that were successfully visited by the robots. It is defined as
    $\rho_{\text{visit}}=\frac{N_{\text{visit}}}{N_{\text{total}}}$
    where $N_{\text{visit}}$ is the number of visited landmarks and $N_{\text{total}}$ is the total number of landmarks. A landmark is considered visited if a robot comes within a certain distance (e.g., 0.1 meters) of the landmark center. The task is regarded as successful if the value of this metric is equal to 1.
  • Pursuing task
    User instruction:Engage in flocking behavior with all robots on the map, moving
    toward the lead robot. The lead robot’s movement is unpredictable, so maintain
    cohesion by staying connected, ensure alignment by moving in sync, and uphold
    separation by keeping a safe personal space. Additionally, be cautious to avoid
    collisions with any obstacles in the environment.
    Evaluation metric: The pursuing task is evaluated based on two metrics. The task is treated as successful when both metrics exceed their corresponding thresholds.
    1) Average distance to prey, denoted as $d_{\text{avg-prey}}$: It measures the average distance between all robots and the prey’s position. It is defined as
    $d_{\text{avg-prey}}=||\mathbf{p}_{\text{avg}}-\mathbf{p}_{\text{prey}}||$
    where $\mathbf{p}_{\text{avg}}$ is the average position of all robots, and $\mathbf{p}_{\text{prey}}$ is the prey’s position. The task is regarded as successful if the value of this metric is less than 1.
    2) Maximum of minimum distances, denoted as $d_{\text{maxmin}}$: It quantifies the largest minimum distance between each robot and its closest neighbor. Its definition is equivalent to the metric used in the aggregation task. The task is regarded as successful if the value of this metric is less than 1.
  • Bridging task
    User instruction:The robots need to evenly form a straight line bridge at the
    position where x is equal to zero within the range of y between minus two and two.
    Evaluation metric: Procrustes distance, denoted as $d_{\text{proc}}$: It quantifies the shape similarity between the robots’ final positions and the target straight line. Its definition is the same as the metric used in the shaping task and hence omitted here. The task is regarded as successful if the value of this metric is less than 0.1.
  • Clustering task
    User instruction:Robots with initial positions in the same quadrant need to cluster in the designated area of that corresponding quadrant.
    Evaluation metric: Achievement Ratio, denoted as $r_{\text{achieve}}$: This metric evaluates the proportion of robots that successfully reach their assigned target regions based on their initial quadrant classification. It is defined as
    $r_{\text{achieve}}=\frac{\sum_{q=1}^{4}N_{q,\text{achieved}}}{N_{\text{total}}}$
    where $N_{q,\text{achieved}}$ represents the number of robots in quadrant $q$ that reach the corresponding target region within a tolerance of $0.1$, and $N_{\text{total}}$ is the total number of robots. The task is considered successful if $r_{\text{achieve}}=1$, indicating all robots meet the criteria.
Details of Software Architecture
We designed a modular architecture consisting of seven core modules, each containing multiple classes (Extended Data Fig. 2). These classes have inheritance, association, and composition relationships, which enhance system design by enabling code reuse, modularity, and flexibility. The Core Module defines the interfaces between modules, ensuring that they can seamlessly integrate into the system as long as they follow these interfaces. The Skill Module handles skill library operations, including the representation of skills as a skill graph and the functionality to construct, modify, and extend this graph. The Action Module contains all action nodes that encompass tasks such as analyzing constraints, designing functions, writing code, performing syntax checks, and debugging, all guided by interactions with the LLM. The Environment Module encompasses various simulation environments or real-world scenarios, the Constraint Module handles constraint-related tasks, the File Module manages file storage, and the Feedback Module processes all feedback.
The core of the architecture is the Core Module, which includes a set of interfaces and base classes that provide shared interfaces and core functionality to the system’s other modules. Specifically, the Core Module uses BaseActionNode, ActionNode, and CompositeActionNode to implement the Composite Pattern[41], ensuring consistent usage of single and composite action nodes, thereby effectively simplifying the system’s complexity. All actions in the Action Module inherit directly from ActionNode, and these action nodes form the core functionality required by the framework. Take GenerateFunctions, a CompositeActionNode, as an example: it consists of four actions—DesignFunctionAsync, WriteFunctionsAsync, GrammarCheckAsync, CodeReviewAsync, and WriteRun—executed in a specific order. This composite node can be reused whenever GenerateFunctions is needed, eliminating the need to rebuild the sequence. Furthermore, GenerateFunctions itself can be treated as a standard ActionNode, maintaining consistency in how single and composite actions are handled.
Moreover, the Core Module provides several key interfaces to support the system’s diverse requirements. The Feedback interface provides a unified handling mechanism for HumanFeedback, CodeBug, and CriticFeedback, as shown in the Feedback Module. The BaseFile interface standardizes the handling of various file types, coverage code files, program logs, and Markdown documents, as shown in the File Module. The BaseEnvironment interface offers standardized access points for different simulation environments, allowing the system to easily adapt to various runtime environments, as shown in the Environment Module. The BaseGraphNode interface unifies the operations of ConstraintNode and SkillNode, ensuring consistency between them and simplifying the establishment of mapping relationships between the two. SkillNode forms SkillLayer, and multiple SkillLayers can form a SkillGraph, constituting the layered structure of the framework mentioned above, as shown in the Skill Module.
The proposed software architecture has the following features. First, by defining generic interfaces and base classes, it achieves a high degree of scalability, allowing the system to easily introduce new functional modules or replace existing ones while maintaining overall system stability. Second, by leveraging the composite pattern technique, which organizes objects into tree-like structures, it unifies the handling of individual and composite skills. Individual skills serve as leaf nodes, while composite skills are represented as branches, allowing users to easily build complex skill structures by combining and nesting different skill nodes. Third, the system supports both simulation and real-world experimental platforms, achieving a unified access point across different platforms.
Details of Automatic Deployment
The following introduces the tools of Ansible and Docker and how they are integrated into our automatic deployment framework.
Ansible is an open-source automation tool that allows tasks to be performed consistently across multiple devices. In our framework, it is used to establish wireless connections with robots via SSH (Secure Shell, enabling secure remote communication) and execute predefined playbooks—scripts that outline the steps for deployment. For example, Ansible ensures directories are created, source code is copied, dependencies are installed, and permissions are set on all robots simultaneously. This consistency reduces human error and eliminates the need for manual intervention on individual robots.
The Docker environment includes all the necessary components for seamless robot operation and code execution. It is equipped with ROS (Robot Operating System), a middleware essential for controlling and managing robotic systems. Additionally, it includes a Python runtime preconfigured with all dependencies required to execute the LLM-generated code.
The deployment process begins with Ansible transferring to each robot the necessary files, such as Python scripts, ROS configuration files, and Dockerfiles, which define the instructions to build the containerized environment for running the code. Once these files are in place, Ansible uses Docker to build the Docker image, packaging the runtime environment and all necessary dependencies. It then pulls and tags prebuilt images to reduce setup time by downloading existing configurations. Afterward, Ansible launches the containers, starting the robot-specific workspace and preparing it for code execution. Inside the container, the code is compiled to ensure compatibility with the ROS environment. Finally, the LLM-generated code is executed via ROS launch files, allowing the experiment to run automatically without further manual intervention.
Data availability
The data in this study are available in the main text and the supplementary information. Other source data are available from the corresponding author upon reasonable request.
Acknowledgments
The authors would like to thank Jialing Lyu for her help in editing the videos. This work was partially supported by the STI 2030-Major Projects (Grant No. 2022ZD0208800). Roderich Groß acknowledges support by the OpenSwarm project, which has received funding from the European Union’s Horizon Europe Framework Program under Grant Agreement No. 101093046.
Author contributions
S.Z., R.G., R.Z., and M.C. designed the research and wrote the paper; W.J., H.C., M.C., G.Z., and L.X. performed research and analyzed data.
Competing interests
The authors declare no competing financial interests.
Correspondence
Correspondence and requests for materials should be addressed to Shiyu Zhao.
Extended Data Fig. 1: Different user instructions for the encircling task and some failure cases. a. Comprehensive user instructions result in a relatively high success rate yet are not guaranteed to succeed. b. Concise user instructions result in a relatively low success rate yet may successfully accomplish the task. c. Overly brief user instructions usually result in a low success rate.
Extended Data Fig. 2: The software class diagram of GenSwarm consists of seven modules. The Core Module defines the interfaces between all modules, ensuring seamless integration. The Skill Module manages the skill library and the construction and modification of skill graphs. The Action Module contains action nodes responsible for tasks. The Environment Module supports various simulation and real-world environments, whereas the Constraint Module handles constraint-related tasks. The File Module manages file storage, and the Feedback Module processes all feedback. These modules interact through standardized interfaces defined by the Core Module, ensuring flexibility and consistency across the system.